OpenStudy (anonymous):

cos2x- sqrt 2 sinx=1 Find all solutions

OpenStudy (campbell_st):

well \[\cos(2x) = 1 - 2\sin^2(x) \] so make the substitution and you'll have a quadratic equation \[1 - 2\sin^2(x) -\sqrt{2}\sin(x)= 1\] rewriting the equation and collecting like terms you get \[\sin^2(x) + \sqrt{2} \sin(x) = 0\] this can now be solved by factoring

OpenStudy (anonymous):

@campbell_st where did the sin^2 come from?

OpenStudy (campbell_st):

ok... quick question is it 1. cos(2x) or 2. \[\cos^2(x)\]

OpenStudy (anonymous):

1. cos2x (there is no exponents)

OpenStudy (campbell_st):

ok... great, the stuff above is correct. cos(2x) can be rewritten is several forms, and they are known facts about cos(2x) to show you if you are familar with the sum of 2 angles cos(x + x) = cos(x) cos(x) - sin(x)sin(x) = cos^2(x) - sin^2(x) (equation 1) now you may know sin^2(x) + cos^2(x) = 1 by making cos^2(x) the subject I get cos^2(x) = 1 - sin^2(x) (equation 2) if I substitute equation 2 into equation 1 you get 1- 2sin^2(x) its just a property you'll need to know.... so, long story short... I made the substitution cos(2x) = 1 - sin^2(x) which means you now have a "quadratic" equation in sin(x) that can be solved for x. sorry to be long winded but I hope it makes sense.

OpenStudy (campbell_st):

oops missing a 2 cos(2x) = 1 - 2sin^2(x)

OpenStudy (anonymous):

umm.. so cosx= pi/3

OpenStudy (campbell_st):

no you have \[2\sin^2(x) + \sqrt{2} \sin(x) = 0\] take sin(x) as out as a common factor and you get \[\sin(x)(2\sin(x) + \sqrt{2}) = 0\] so you are looking where does sin(x) = 0 and where does \[2\sin(2) + \sqrt{2} = 0\] these will be the solutions...

OpenStudy (anonymous):

so sinx= 0 +2pi..etc and sinx=5pi/4+2pi

OpenStudy (campbell_st):

ummm \[\sin(x) = -\frac{\sqrt{2}}{2} \] 3rd quadrant is \[\frac{5\pi}{4}\] and 4th quadrant \[\frac{7\pi}{4}\]

OpenStudy (campbell_st):

hope that makes sense...

OpenStudy (anonymous):

yes its does i get that but has to be in terms or radians

OpenStudy (campbell_st):

you just need the 3rd angle in general solution form

OpenStudy (anonymous):

ok i think i got it, thanks.

OpenStudy (campbell_st):

well you can either use \[\frac{7\pi}{4} + 2\pi\]

OpenStudy (anonymous):

yeah i just put all of them